### v2

# Vision-Encoders (Already) Know What They See: Mitigating Object Hallucination via Simple Fine-Grained CLIPScore

Abstract
Recently, Large Vision-Language Models (LVLMs) show remarkable performance across various domains. However, these models suffer from object hallucination. This study revisits the previous claim that the primary cause of such hallucination lies in the limited representational capacity of the vision encoder. Our analysis reveals that the capacity of the vision encoder itself is already adequate for detecting object hallucination. Based on this insight, we propose a Fine-grained CLIPScore (F-CLIPScore), a simple yet effective evaluation metric that enhances object-level granularity by incorporating text embeddings at the noun level. Evaluations on the OHD-Caps benchmark show that F-CLIPScore significantly outperforms conventional CLIPScore in accuracy by a large margin of 39.6% without additional training. We further demonstrate that F-CLIPScore-based data filtering reduces object hallucination in LVLMs (4.9% in POPE).
Vision-Encoders (Already) Know What They See:
Mitigating Object Hallucination via Simple Fine-Grained CLIPScore
Hongseok Oh  Wonseok Hwang University of Seoul {cxv0519, wonseok.hwang}@uos.ac.kr
1 Introduction
CLIPScore (w/o training): A lady and two children in the street playing with a tennis racquet, a car nearby, and a chair.
CLIPScore (trained): A lady and two dogs in the park playing with a frisbee.
F-CLIPScore (w/o training): A lady and two children in the street playing with a tennis racquet.
Figure 1: A representative example from the OHD-Caps test set is shown. The original CLIP selects a sentence mentioning “children” and “tennis” but adds hallucinated objects. The OHD-Caps-trained CLIP hallucinates “dog” and “frisbee” without introducing new content. In contrast, F-CLIPScore selects a sentence that preserves the original meaning without hallucinations.
Recent studies identify Large Vision-Language Models (LVLMs) as a leading approach for vision-language integration Liu et al. (2023); Wang et al. (2024a); Zhu et al. (2023); Chen et al. (2024). However, like hallucinations in large language models (LLMs) Ji et al. (2023); Zhang et al. (2023), LVLMs also exhibit object hallucination—describing nonexistent or misidentified objects—which undermines their reliability Li et al. (2023a); Liu et al. (2024b).
Liu et al. (2024c) built OHD-Caps, a dataset designed to measure object hallucination. The dataset comprises 1.5k image and captions pairs where a model needs to select the best caption that does not show hallucinations. They found the CLIPScore shows the accuracy 10–20% and the further fine-tuning with the proposed objective function with the dataset enhance the accuracy up to 80–90%. However, when they connected the OHD-Caps-trained CLIP Radford et al. (2021) to an LVLM and conducted full fine-tuning, the resulting accuracy sometimes drops, showing lower performance compared to original CLIP (for instance, 1st row of Table 4 in (Liu et al., 2024c) shows the accuracy in POPE benchmarks (Li et al., 2023b) drops from 85.4% to 81.2%).
Also, as shown in Figure 1, OHD-Caps-trained CLIP tends to hallucinate by replacing existing objects. When both CLIPScore and OHD-Caps-trained CLIPScore fail but F-CLIPScore succeeds, a clear pattern emerges: CLIPScore adds nonexistent objects, while OHD-Caps trained CLIPScore replaces existing ones. This trend holds across COCO, Flickr30k, and NoCaps with rates of 56%, 58%, and 59%, respectively. This suggests that object hallucination may stem from factors beyond the vision encoder’s capacity.
To address this issue, we introduce Fine-grained CLIPScore (F-CLIPScore), a novel image-text correlation metric. F-CLIPScore leverages a sentence parser like spaCy Honnibal et al. (2020) and the forward pass of a Vision-Language Model (VLM) like CLIP Radford et al. (2021), offering an efficient way to evaluate the vision encoder’s representational capacity. Our experimental results show that applying F-CLIPScore to the OHD-Caps test set improves accuracy by +39.6% without additional training. This indicates that the limited capacity of the vision encoder may not be the primary cause of object hallucination. Additionally, we verify that using F-CLIPScore for pretraining data curation in LVLMs enables the training of models with reduced hallucination, even with significantly fewer data samples. Notably, in LVLM pretraining, data filtering alone improved POPE accuracy by 4.9% compared to the baseline.
This study offers the following key contributions:
  • We introduce Fine-grained CLIPScore, a novel evaluation metric that relies solely on forward propagation.
  • We provide evidence that the object hallucination does not actually primarily stem from the vision encoder’s capacity.
  • Demonstrating that F-CLIPScore enables more efficient LVLM training with reduced object hallucination through pretraining data curation.
Our code is available at https://github.com/abzb1/f-clip
2 Related Work
Object hallucination refers to cases in which the generated textual descriptions include objects that do not correspond to the given image Liu et al. (2024b). LVLMs generally consist of three components: a vision encoder, an LLM, and an adapter Liu et al. (2023). The structural characteristics of LVLMs contribute to object hallucination, which arises from multiple intertwined factors Liu et al. (2024b). While some studies argue that hallucinations can be mitigated by enhancing the decoding process of the LLM Manevich and Tsarfaty (2024); Wang et al. (2024b), others suggest that the root cause lies in the limited representational capacity of the vision encoder Liu et al. (2024c). Additionally, some research indicates that training the adapter with contrastive data is essential to reduce object hallucination Jiang et al. (2024). And the trained bias of the model has also been identified as a cause of hallucination Hu et al. (2023); Liu et al. (2024a).
CLIPScore Hessel et al. (2021) is a reference-free evaluation method that assesses the consistency between an image and text caption by computing the cosine similarity between the embeddings generated by the vision encoder and text encoder of the CLIP model. Beyond its application in measuring caption quality, several studies have also leveraged CLIPScore for data curation in the training of Vision-Language Models (VLMs) Schuhmann et al. (2021); Gadre et al. (2023).
A recent study Liu et al. (2024c) utilized CLIPScore to evaluate object hallucination in Vision-Language Models (VLMs). Their findings suggest that this phenomenon stems from the limited capacity of the vision encoder. In this study, we carefully reassess this claim and demonstrate that object hallucination is not necessarily caused by the limitations of the vision encoder alone.
3 Methods
3.1 Motivation
Figure 2: Histograms of cosine similarity between two embedding vectors: one from the original CLIP-L and the other from the OHD-Caps-trained CLIP-L. (a) The histogram from the vision encoders. Correct (blue) indicates the scores are from the examples where original CLIP-L predict the ground truth. The other examples are colored in orange. (b) The histogram from the text encoders. Same color scheme is employed. Measured only on ground truth text. (c) The cosine similarity distribution between text embeddings of text without object hallucination (purple) and with object hallucination text (green) for all samples.
Based on our initial observation that OHD-Caps-trained CLIP does not necessarily yield better results (Section 1), we further investigate how fine-tuning affects the embedding vectors produced by vision and text encoders of CLIP. We compute the cosine similarity between two embedding vectors: one from the original CLIP-L and the other from the OHD-Caps-trained CLIP-L, using image-text pairs in the OHD-Caps test set.
We first computed the distribution of cosine similarity between the image embeddings from CLIP-L and OHD-Caps-trained CLIP-L for the samples that were correctly predicted by CLIP-L. We then performed the same analysis on the samples that were incorrectly predicted by CLIP-L and compared the two distributions. As shown in Figure 2a, we observe that there is no substantial difference between the two distributions. We conducted the same analysis using ground-truth text embeddings and similarly found no significant difference in the cosine similarity distributions between the pre-trained and fine-tuned models for samples that were correctly and incorrectly predicted by CLIP-L Figure 2b. In contrast, Figure 2c reveals significant changes in text embeddings for captions with hallucination (purple), and without hallucination (green) highlighting the distinct adaptation of text representation. These observations suggest that OHD-Caps training may shift the representation space of the vision encoder while enhancing the text encoder’s discriminative ability.
3.2 Fine-grained CLIPScore
Motivated by above observations, we propose a simple metric called Fine-grained CLIPScore (F-CLIPScore), which enhances the discriminative power of the VLM in order to utilize textual information with more granularity without additional training. F-CLIPScore first utilizes the spaCy parser Honnibal et al. (2020) to extract nouns from a given sentence. Then, it evaluates the quality of an image caption by averaging the CLIPScore of the entire sentence and each individual noun (Figure 3). Mathematically, given a full sentence $S$ and a total of $N$ nouns, denoted as $n_{i}$, F-CLIPScore is defined as
$\operatorname{F-CLIPScore}(S)=\frac{\operatorname{CLIPScore}(S)+\sum_{i=1}^{N}{\operatorname{CLIPScore}(n_{i})}}{N+1}$ (1)
Figure 3: The graphical representation of F-CLIPScore.
4 Experiments
For evaluating OHD-Caps test set, F-CLIPScore only need extra parser (we used en_core_web_sm) time (less than a second per sentence). For training CLIP with equation 2 on the OHD-Caps train set, we used an batch size of 64 and a learning rate of 1e-5, which required 3 hours on an H100 GPU. For LLaVA pretraining, we employed an effective batch size of 128 and a learning rate of 1e-3, which took 7 hours on two H100 GPUs.
5 Results
By utilizing the F-CLIPScore, which directly leverages the image embeddings from the CLIP vision encoder without any training or gradients, while only adding a parsing process during forward propagation, we can efficiently gain insights into whether the issue of object hallucination arises from the limited capability of the vision encoder.
5.1 F-CLIPScore on OHD-Caps
We evaluated the OHD-Caps Liu et al. (2024c) test set, an object hallucination assessment dataset, using the proposed F-CLIPScore. As shown in Table 1, evaluation results with CLIP ViT-L Radford et al. (2021) indicate that F-CLIPScore outperformed the baseline model by up to 39.6% without additional training (row 4 vs. 5). However, it still performed 21.8% to 38.6% worse than trained models (row 5 vs. row 7). This may indicate that although CLIP vision encoders may not be the main cause of object hallucination, further training may enhance their capability.
OHD-Caps ACC (%, $\uparrow$)
Metric coco flickr30k nocaps
w/o training
CLIPScore 22.6 22.6 12.4
$\operatorname{F-CLIPScore}$ 62.2 62.2 46.6
trained w/ OHD-loss
CLIPScore 79.8 ±1.7 84.8 ±1.6 84.0 ±0.7
trained w/ F-CLIPScore loss
CLIPScore 80.5 ±1.8 84.8 ±1.3 84.1 ±1.3
$\dagger$: We trained openai/clip-vit-large-patch14 with the train code from Liu et al. (2024c)
Table 1: Accuracy on the OHD-Caps test set evaluated with OpenAI CLIP-L. For trained models, we report the average evaluation results over 10 runs with different random seeds.
To test whether F-CLIPScore is orthogonal to the previously proposed method Liu et al. (2024c), we modify the loss as
$L_{CLIP}+L_{OHD}+\frac{\alpha}{B}\sum_{i=1}^{B}{(1-\operatorname{F-CLIPScore}(I_{i},C_{i}))}$ (2)
, where $L_{CLIP}$ is the contrastive loss proposed in  Radford et al. (2021), and $L_{OHD}$ is the marginal loss proposed in  Liu et al. (2024c). $B$ denotes the batch size, while $I_{i}$ and $C_{i}$ are the $i$-th positive pair. $\alpha$ is a hyperparmeter that we set to $0.3$. Applying F-CLIPScore as a loss improved performance by 0.7 and 0.1 percentage points on the COCO and Flickr30K subsets, respectively (row 7 vs. 9).
5.2 LVLM Pretrain Data Curation with F-CLIPScore
As shown in Section 5.1, F-CLIPScore was able to exhibit competent performance in detecting object hallucination without any training, and could further serve as an orthogonal method for training VLMs.. We aimed to investigate whether F-CLIPScore can influence the pretraining process of LVLM by acting as a data filter between the vision encoder and the LLM. To this end, we applied F-CLIPScore to filter the pretraining data for LLaVA Liu et al. (2023), which consists of 558k samples, by removing the bottom $x\%$ of the alignment training set based on F-CLIPScore.
POPE Acc ($\uparrow$, %) Filtering rate (%)
Filtering Method 20 30 40 50 60
Random 50.7 49.9 49.9 50.5 49.8
CLIPScore (w/o train) 52.2 47.3 50.0 53.1 50.0
F-CLIPScore (w/o train) 51.6 55.5 50.6 50.1 51.8
CLIPScore (trained) 49.8 49.8 50.8 49.8 52.6
w/o filtering 50.58
Table 2: POPE benchmark accuracy after LLaVA pretraining. We use CLIP-L Radford et al. (2021) as a vision encoder, and Llama 2 7B Touvron et al. (2023) as an LLM backbone. “Trained” indicates the OHD-Caps-trained CLIP-L. “Random” refers to the removal of $x\%$ of samples chosen at random.
As shown in Table 2, training the alignment model on the top 70% of data curated by F-CLIPScore resulted in a +4.9% improvement in POPE accuracy compared to training on the entire dataset (second row). In contrast, using the OHD-Caps-trained CLIP for filtering yielded only marginal gains, and random filtering showed no improvement. These results suggest that F-CLIPScore effectively captures object hallucination-related quality, even in general-purpose datasets. This finding underscores the need to explore alternative causes of object hallucination beyond the capacity of the vision encoder. Although model performance shows a non-linear trend with respect to the filtering ratio, our experiments used a fixed number of training epochs, which may account for this behavior. Goyal et al. (2024) highlights the need to balance data quality with computational cost, and our findings align with this perspective. Identifying the optimal filtering ratio under such trade-offs remains an important direction for future research.
6 Conclusion
We introduce F-CLIPScore, a simple yet effective metric for evaluating fine-grained image-caption alignment and addressing object hallucination in Vision-Language Models. Unlike conventional CLIPScore, which relies solely on sentence-level embeddings, F-CLIPScore also incorporates noun-level embeddings. This refinement allows the model to better mitigate object hallucination without requiring additional training for the vision encoder. We validate F-CLIPScore by showing a +39.6% accuracy in OHD-Caps benchmark. We also show that data filtering based on F-CLIPScore can enhance LVLM performance in hallucination mitigation, even with a reduced dataset. These results may indicate that the limitations of existing evaluation metrics are one possible origin of object hallucination, which fails to efficiently reveal the capacity of the vision encoder.
Limitations
While this study proposes a method for analyzing and mitigating object hallucination using F-CLIPScore, it is subject to the following limitations. First, we were unable to conduct experiments on the Supervised Fine-Tuning (SFT) for Large Vision-Language Models (LVLMs). In the LVLM training pipeline, after the alignment pretraining phase—where the vision encoder and LLM remain frozen—the SFT stage follows, in which these components are unfrozen and further trained. However, due to computational resource limitations, we did not fully explore the potential impact of F-CLIPScore during SFT. Future work could investigate ways to incorporate F-CLIPScore into the SFT process to enhance training effectiveness.
Second, our method faces linguistic constraints and challenges in multilingual generalization. This study employs the spaCy parser Honnibal et al. (2020) to extract nouns from text, a technique that performs relatively reliably in well-structured languages such as English. However, parsing accuracy may vary across different languages, potentially leading to inconsistencies in F-CLIPScore computation. To address this, future research should explore the scalability of F-CLIPScore by evaluating its effectiveness on multilingual datasets and refining the parsing methodology for broader linguistic applicability.
Appendix A Additional Results on F-CLIPScore Robustness
We randomly injected nouns extracted from a Wikipedia corpus Davies (2015) into the caption embeddings to examine how F-CLIPScore responds to such perturbations. As shown in Table 3, we observed the expected decline in accuracy as nouns were replaced. This indicates that the strong performance of F-CLIPScore, even without additional training, is not due to random noise, but rather stems from its ability to more effectively leverage fine-grained textual embeddings.
Injection Rate
0.2 0.4 0.6 0.8 1.0
COCO Acc mean $\pm$ std (%, $\uparrow$)
38.9$\pm$2.5 35.9$\pm$1.4 31.0$\pm$1.7 23.5$\pm$2.0 19.9$\pm$1.3
Flickr30k Acc mean $\pm$ std (%, $\uparrow$)
37.5$\pm$1.5 33.0$\pm$1.2 26.2$\pm$2.1 22.0$\pm$1.6 18.9$\pm$3.0
NoCaps Acc mean $\pm$ std (%, $\uparrow$)
29.4$\pm$1.9 24.2$\pm$2.3 21.2$\pm$2.0 16.2$\pm$1.7 16.5$\pm$0.4
Table 3: F-CLIPScore results on the OHD-Caps test set with different random noun injection rates (0.2 to 1.0). Each value shows the mean accuracy $\pm$ standard deviation over five random seeds.
OHD-Caps ACC (%, $\uparrow$)
Metric coco flickr30k nocaps
CLIPScore 22.6 22.6 12.4
$\operatorname{F-CLIPScore}_{V}$ 23.6 24.8 15.8
$\operatorname{F-CLIPScore_{NP}}$ 54.2 57.8 39.6
$\operatorname{F-CLIPScore}_{N}$ 62.2 62.2 46.6
Table 4: Accuracy on the OHD-Caps test set evaluated with OpenAI CLIP-L. $\operatorname{F-CLIPScore}_{N}$ refers to the noun-based $\operatorname{F-CLIPScore}$ defined in Equation 1. $\operatorname{F-CLIPScore_{NP}}$ replaces the nouns with noun phrases, and $\operatorname{F-CLIPScore}_{V}$ replaces the nouns with verbs.
We experimented with alternative configurations of F-CLIPScore using verbs and noun phrases instead of nouns. As shown in Table 4, using verbs yielded little to no improvement, while noun phrases performed better than verbs but still fell short of the performance achieved with nouns.
Appendix B Licenses for the datasets and models used
LLaMA 2 7B is released under the LLaMA Community License. The LLaVA pretraining dataset is licensed under a combination of LAION, CC, and SBU licenses. The OpenAI CLIP model is distributed under the MIT License. COCO is available under the Creative Commons Attribution 4.0 License (CC BY 4.0). Flickr30k is provided under a custom license that permits use for non-commercial research and/or educational purposes. NoCaps is distributed under the Creative Commons Attribution 2.0 License (CC BY 2.0). We used the datasets and models in accordance with their respective licenses.